OpenStudy (anonymous):

PB has endpoints at -2 and 12. What is the coordinate of its midpoint? Ps. PB has a line on top of the letters. gladly love the help

OpenStudy (mathstudent55):

All you need to do is find the average of -2 and 12. Do you know how to find the average of two numbers?

OpenStudy (anonymous):

add the two numbers, and divide their sum by two

OpenStudy (anonymous):

i think?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

Yes, that's it.

OpenStudy (anonymous):

soo -2+12=10

OpenStudy (mathstudent55):

Right.

OpenStudy (anonymous):

and 12/2=6?

OpenStudy (mathstudent55):

No. You use the sum. Since -2 + 12 = 10, now do 10/2, not 12/2.

OpenStudy (anonymous):

5

OpenStudy (mathstudent55):

That's it.

OpenStudy (anonymous):

So what do i write for my answer. just put 10 and 5?

OpenStudy (mathstudent55):

Now we can check. If 5 is the midpoint between -2 and 12, it has to be the same distance from both -2 and 12. |5 - (-2)| = |7| = 7 |5 - 12| = |-7| = 7 Yes, 5 is exactly 7 units from both -2 and 12, so 7 is the midpoint. The answer is only 5.

OpenStudy (anonymous):

THANK U SO MUCH< U HELPED ME UNDERSTAND ITS SO EASY I MAKE IT LOOK HARD

OpenStudy (mathstudent55):

You are very welcome.